OpenStudy (anonymous):

what's arccsc(-2) equal to

jimthompson5910 (jim_thompson5910):

Let x = arccsc(-2) x = arccsc(-2) csc(x) = csc(arccsc(-2)) csc(x) = -2 1/sin(x) = -2 1 = -2*sin(x) 1/(-2) = sin(x) -1/2 = sin(x) sin(x) = -1/2 Now use the unit circle to find that sin(-pi/6) = -1/2 So x = -pi/6 This means that arccsc(-2) = -pi/6

OpenStudy (anonymous):

so it's -30 degrees

jimthompson5910 (jim_thompson5910):

yes, since -pi/6 = -30 degrees

OpenStudy (anonymous):

will it be 330 degrees or 30 degrees? it's a multiple choice question and 330 and 30 are both the answers

jimthompson5910 (jim_thompson5910):

Well 330 is a coterminal angle to -30 degrees

OpenStudy (anonymous):

so would you choose 330 or -30

jimthompson5910 (jim_thompson5910):

I would go with -30 since that's closest to 0

OpenStudy (anonymous):

yay. you're absolutely right thank you very much!